find the slope of the line having the given slope and containing the given point m=4/5,(8,-8)
Answer by stanbon(75887)

find the slope of the line having the given slope and containing the given point m=4/5,(8,-8)
---
The slope is 4/5.
I assume you want the equation:
---
Equation Form: y = mx+b
Substitute and solve for "b":
-8 = (4/5)(8)+b
-8 = (32/5) + b
---
-40/5 = (32/5) = b
b = -72/5
----
Equation:
y = (4/5)x-(72/5)
===================
Cheers,
Stan H.
==========